OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 4x – 5y = –18 3x + 2y = –2 A. (3, –1) B. (–2, 7) C. (–1, 3) D. (–2, 2) Is it D?

OpenStudy (phi):

multiply the first equation by 2 multiply the second equation by 5 you get these two equations, which you add together: 8x – 10y = –36 15x + 10y = –10 --------------- 23x + 0 = -46 23x= -46 divide both sides by 23 to get x = -46/23 = -2

OpenStudy (phi):

to find y, we can use either of the original equation. The 2nd equation has smaller numbers, so let's use that one 3x + 2y = –2 replace x with -2: 3*-2 + 2y = -2 -6 + 2y = -2 add +6 to both sides -6+6 + 2y = 6-2 2y = 4 y=2 so (-2,2) is the answer. we can check that this works in the first equation if we want to make sure we did not make a mistake.

OpenStudy (anonymous):

Thank you! :D

OpenStudy (phi):

yw